Cuzziculum Vitae

Name Vanq and Margo

Date of Birth 15 June 1983

Plase of Birth Plovdiv,Bulgaria

Nationality Bulgarian

Marital Ststus singal

Telefone Number +359-988-248458 (phone) +359-32-234-479 (work)

E-mai gtd@avb.bg Education 1993-1998 Masters Degree in Biology \underline{University of Plovdiv}

Average grade: Excellent (5.53) 1988-1992 Mathematical high Scoool of Plovdiv

Average Grade; Excellent (5.87)

Current Occupation Imformation specialist at Biochem LTD